# Image is All You Need: Towards Efficient and Effective Large Language Model-Based Recommender Systems

### v1

@@ 4.3.3. Analysis on Robustness
Figure 7. Performance across item groups, where higher IDs indicate warmer item groups.
4.4. Ablation Studies (RQ3)
In Table 2, we conduct ablation studies to understand the effectiveness of each component in I-LLMRec. The variant without any additional components is trained solely using Equation 5 and recommends items based only on image features (row (a)). We make the following observations: 1) Effect of ILA: We observe that equipment with the ILA module significantly improves performance (row (a) vs. row (b)). This indicates that the ILA module effectively aligns visual features with the language space, enhancing the LLM’s ability to understand user preferences. 2) Effect of extending to multiple feature types in IRE: We observe that incorporating CF features along with image features [^10] The inclusion of specific feature types is used for both the model training in Equation 7 and during inference in Equation 8. yields better performance than using either the image or CF feature alone (row (b), (c) vs. (d)). Moreover, integrating all three feature types (i.e., Image, CF, and Text) further improves performance (row (d) vs. (e)). This demonstrates that extending multiple feature types enhances recommendation effectiveness beyond image features alone.
Effectiveness of Feature types across cold/warm items. To further analyze the impact of different feature types, we evaluate their performance on cold and warm item recommendations. Specifically, we follow (Liu et al., 2023) by sorting the items in ascending order based on the frequency of interaction and dividing them into five equally sized groups, where a higher group ID indicates a warmer item group, Then, we compare the recommendation performance across these groups. As shown in Figure 7, we observe that using image features (blue bar) generally outperforms CF features (orange bar) for cold items (lower ID groups). However, as the item group transitions from cold to warm, CF’s effectiveness improves, eventually surpassing Image in Item Group 5 (Warm). This suggests that CF is particularly effective for recommending warm items. With these insights, combining Image and CF (green bar) helps mitigate CF’s weakness in recommending cold items, resulting in higher performance than CF alone in Item Group 1-2. At the same time, the CF features compensate for weakness of the image features in recommending warm items, leading to better performance than image alone in Item Group 4-5. Furthermore, adding textual features (red bar) further boosts performance across all Item Groups, exceeding the performance of Image+CF. This analysis highlights how different feature types complement each other to enhance recommendation.
Table 3. Performance when item images are missing.
@@ 4.4. Ablation Studies (RQ3)
4.5. Missing Image Scenario (RQ4)
@@
In this section, we explore how to handle scenarios where item images are missing and evaluate how effectively we can manage them. Specifically, we randomly remove 50% of the item images from the entire item pool. As a result, these missing images cannot be used for item semantics representation (Equation 2) and retrieval-based recommendation (Equation 5 and Equation 8). To compensate for the missing images, we substitute them with readily available textual descriptions. Similarly, for retrieval-based recommendations, when item images are missing, we extract textual features derived from descriptions using SigLIP (Zhai et al., 2023) text encoder rather than extracting visual features from images using SigLIP visual encoder. As shown in Table 3, I-LLMRec trained with half of the items missing images outperforms the baselines (TALLRec and TRSR), even though the baselines were trained without any missing item images. This demonstrate the effectiveness of I-LLMRec in handling the missing image scenario, which shows the practicality of I-LLMRec in reality.
5. Related Work
LLM-based Recommendation. Due to the power of textual understanding and world knowledge of LLMs, a myriad of studies have spurred the development of LLMs for recommender systems (Bao et al., 2023; Wang and Lim, 2023; Kim et al., 2024a; Zhang et al., 2023; Liu et al., 2024a). To adapt LLMs for recommendation tasks, previous studies represent the interaction history as a sequence of item titles. Additionally, some studies enhance recommendation performance by incorporating item attributes, providing richer item semantics to the LLMs (Attribute-based Representation). Specifically, TALLRec (Bao et al., 2023) converts item IDs into their titles and captures user preference via LoRA (Hu et al., 2021) fine-tuning. TransRec (Lin et al., 2024) improves the user preference understanding by including item titles and attributes in prompts. IDGenRec (Tan et al., 2024) incorporates item titles generated by an independent LLM for unique and meaningful semantics. More recently, to feed the richer item semantics to LLMs, a description of items have been utilized (Description-based Representation). Specifically, TRSR (Zheng et al., 2024) utilizes a large LLM (i.e., LLaMA-30B-instruct (Touvron et al., 2023a)) to summarize item descriptions, and feeds them to a smaller LLM (i.e., LLaMA-2 7B (Touvron et al., 2023b)) to capture the condensed meaningful information, while ONCE (Liu et al., 2024a) adopts GPT-3.5 (Brown et al., 2020) for the summarizing task. Despite their notable success to capture the richer item semantics, they face a trade-off between efficiency and effectiveness when representing items in natural language, leading us to utilize image information to address this trade-off. Meanwhile, VIP5 (Geng et al., 2023) and UniMP (Wei et al., 2024) have attempted to incorporate image information into LLMs. However, they lack an in-depth investigation into the overlap between image-description pairs and fail to provide effective alignment strategies between image and language spaces tailored to the recommendation context.
6. Conclusion
In this work, we address the trade-off between efficiency and effectiveness when representing item semantics in natural language for an LLM-based recommender system. Based on the observation that there exists a significant information overlap between images and descriptions associated with items, we propose I-LLMRec, a novel method that leverages images to capture rich item semantics of lengthy descriptions with only a few tokens, thereby capturing user preferences efficiently and effectively. However, as the image and language space are not inherently aligned, we propose the Image-LLM Alignment (ILA) mode, which effectively bridges the gap between these spaces. Furthermore, we propose the Image-based REtrieval (IRE) mode, a retrieval-based recommendation approach, to enhance both the reliability and efficiency of the recommender systems. Our extensive experiments demonstrate that I-LLMRec outperforms natural language-based approaches in terms of both efficiency and effectiveness.
Appendix A Ethics Statement
To the best of our knowledge, we comply with the KDD Code of Ethics and do not violate any ethical issues. Throughout this paper, we use publicly available data and code.
Figure 8. Prompt for summarizing descriptions. The red-colored text denotes the actual metadata for the item.
Appendix B Prompt for Summarization
Figure 8 shows the prompt used for description summarization. Following TRSR (Zheng et al., 2024), we incorporate attribute information along with the description to generate a concise summary.
Appendix C Recommendation Protocol
In this section, we describe a comprehensive explanation of the recommendation protocol, which is consistently applied to Attribute-based Representation, Description-based Representation (i.e., Full Description and Summarized Description), and I-LLMRec as introduced in Section 1. Note that this recommendation protocol follows a retrieval-based recommendation approach similar to the IRE module discussed in Section 3.3. Therefore, we recommend reading Section 3.3 first for a deeper understanding. At a high level, we modify the way items are represented by replacing the visual representation with either an attribute-based or description-based representation within Equation 2, followed by a recommending process similar to IRE module.
[…]
Figure 9. Cosine similarity between item image-description pairs in Amazon Phone and Grocery datasets and image-caption pairs in the COCO dataset using CLIP (Radford et al., 2021).
Figure 10. Sigmoid-based similarity (ranging from 0 to 1) between item image-description pairs across four datasets (Amazon Sport, Art, Grocery, and Phone datasets) using SigLIP (Zhai et al., 2023).
Appendix D Consistent Observation of Information Overlap
We further investigate the information overlap between image and description associated with items, which was explored in Section 1, by expanding these experiments to additional datasets and leveraging another CLIP=based model. 1) Expanding to additional datasets: Building upon our previous comparative experiments in Amazon Sport and Art—where we used CLIP (Radford et al., 2021) to measure cosine similarity in Section 1—we extend our analysis to the Amazon Grocery and Phone datasets. As shown in Figure 9, the average similarity for Amazon Phone and Grocery is approximately 0.31, which is higher than 0.26 average similarity observed in well-aligned COCO image-caption pairs. This consistent observation across multiple datasets further supports the information overlap between item images and descriptions. 2) Leveraging another CLIP model: To further explore the information overlap using a different CLIP variant, we select SigLIP (Zhai et al., 2023), a variant of CLIP with sigmoid loss function, to compute the similarity scores ranging from 0 to 1. Specifically, when we apply SigLIP to measure similarity between item image-description pairs across all four datasets (Amazon Sport, Art, Grocery, and Phone), we observed that, as show in Figure 10, the majority of items exhibit similarity scores close to 1, further demonstrating the significant information overlap between their image and description pairs.
Table 4. Performance comparison when LLM backbone of  I-LLMRec is fine-tuned using LoRA (Hu et al., 2021).
Sport Art
@@ Appendix D Consistent Observation of Information Overlap
LLM Frozen 0.4570 0.3711 0.5883 0.4839
LLM fine-tuning with LoRA 0.4633 0.3772 0.5958 0.4949
Appendix E Frozen LLM vs LoRA Fine-Tuned LLM
Note that fine-tuning the LLM backbone is not mandatory for I-LLMRec. Nevertheless, in this section, we explore potential performance gains from fine-tuning the LLM, a common approach in existing LLM-based recommender systems (Lin et al., 2024; Bao et al., 2023; Tan et al., 2024; Zheng et al., 2024; Liu et al., 2024a). However, since fully fine-tuning an LLM demands substantial computational resources, we adopt the parameter-efficient fine-tuning strategy, LoRA (Hu et al., 2021), which introduces a small number of learnable rank matrices into each LLM’s transformer layer. As shown in Table 4, we observed the performance improvement through fine-tuning, indicating that there is room for further improvement within I-LLMRec. However, considering the computational cost, we opt to freeze the LLM to preserve its intrinsic knowledge while improving training efficiency.
Table 5. Data statistics after preprocessing.
Dataset # Users # Items # Interactions Avg. Len
Sport 24,665 9,884 179,491 7.28
Grocery 84,896 25,632 739,628 8.71
Art 14,986 5,801 121,917 8.13
Phone 59,307 19,671 403,194 6.80
Appendix F Data Statistics
Table 6. Impact of the different backbones.
@@ Appendix F Data Statistics
Appendix G Impact of Different LLMs and Vision Encoders
@@
To investigate the impact of the backbone models on the recommendation performance of  I-LLMRec, we evaluated our model by replacing different backbone models. For the LLM, we used Sheared LLaMA-2.7B (Xia et al., 2023) and Vicuna 1.5-7B (Chiang et al., 2023), an improved version of LLaMA-7B (Touvron et al., 2023a). For the vision encoder, we used CLIP [^11] https://huggingface.co/openai/clip-vit-large-patch14-336 (Radford et al., 2021) and SigLIP [^12] https://huggingface.co/google/siglip-so400m-patch14-384 (Zhai et al., 2023), with the advanced version of CLIP scaling with sigmoid loss. Table 6 shows the following observations: 1) Given the same vision encoder, I-LLMRec with Vicuna 1.5-7B outperforms the model with LLaMA-2.7B. This improvement is likely attributed to Vicuna 1.5-7B’s superior semantic reasoning capability, allowing it to capture user preferences more effectively. 2) Given the same LLM backbone, I-LLMRec with SigLIP performs better than the version with CLIP. This suggests that models that better capture visual features are more effective in delivering item semantics, thereby enabling the LLM to better understand the item semantics. In summary, employing more powerful LLMs and vision encoders leads to improvements in recommendation performance, as these models effectively capture item semantics and user preference.
Appendix H Training Strategy
Throughout this paper, we optimize our model using an end-to-end training strategy to ensure ease of implementation and learning efficiency. However, in this section, we explore the impact of an alternative two-stage training strategy, where we separate the ILA and IRE modules. Specifically, for stage 1, we train the adaptor $M$ through the ILA module to initially align the visual space with the language space. For stage 2, we freeze the adaptor and train only $F_{u}^{*}$ and $F_{i}^{*}$ (*=[Img, CF, Text]) through the IRE module. As shown in Table 7, the two-stage strategy shows performance comparable to the end-to-end strategy, indicating that neither strategy offers a clear advantage over the other. Given this, we opt for the end-to-end training strategy to facilitate the ease of implementation and training efficiency.
Table 7. Training Strategy.
Training Strategy Sport Art
Hit@5 NDCG@5 Hit@5 NDCG@5
End-to-End 0.4570 0.3711 0.5883 0.4839
Stage1+Stage2 0.4604 0.3747 0.5858 0.4887
Figure 11. Prompt template for ILA module. The red-colored text denotes the actual information of an item.
Appendix I ILA Prompt Templates
Figure 11 shows the prompt template for ILA module of  I-LLMRec. As described in Section 3.2, we utilize the prompt templates to align the visual feature with the language space of LLMs by training the adaptor network $M$.

### v2

@@ E.6 Effectiveness of Image Features on Unseen Items
Figure 11: Performance across user groups, where higher IDs indicate warmer user groups.
E.7 Effectiveness of Feature Types Across Cold/Warm Users
We examine how different feature types affect performance under both cold- and warm-user recommendation settings. Following the same protocol as in Appendix 4.4 of the main paper, we partition users into groups based on their sequence length $|\mathcal{S}_{u}|$, ensuring that each group contains the same number of users. As shown in Figure 11, image features (blue bars) consistently outperform CF features across all user groups, indicating that image features remain effective regardless of the user’s interaction history length. Moreover, incorporating more feature types—progressing from Image+CF (green bar) to Image+CF+Text (red bar)—yields steady performance gains in every group, demonstrating that leveraging the full set of features is broadly beneficial. Interestingly, we observe a gradual performance drop across all feature types as user sequence length increases, which contrasts with trends commonly observed in traditional CF models (Kim et al., 2023). We hypothesize that this decline arises from the difficulty LLMs face in capturing user preferences when longer histories introduce greater diversity in the categories of interacted items. Addressing this challenge—accurately modeling user preferences as interaction histories grow—remains an important direction for future work.
Table 8: Performance when item images are missing.
@@ E.7 Effectiveness of Feature Types Across Cold/Warm Users
E.8 Missing Image Scenario
@@
In this section, we explore how to handle scenarios where item images are missing and evaluate how effectively we can manage them. Specifically, we randomly remove 50% of the item images from the entire item pool. As a result, these missing images cannot be used for item semantics representation ($\mathbf{P}_{i}$) and retrieval-based recommendation (Equation 1 and $rec_{u}^{k}$ in the main paper). To compensate for the missing images, we substitute them with readily available textual descriptions. Similarly, for retrieval-based recommendations, when item images are missing, we extract textual features derived from descriptions using SigLIP (Zhai et al., 2023) text encoder rather than extracting visual features from images using SigLIP visual encoder. As shown in Table 8, I-LLMRec trained with half of the items missing images outperforms the baselines (TALLRec and TRSR), even though the baselines were trained without any missing item images. This demonstrate the effectiveness of I-LLMRec in handling the missing image scenario, which shows the practicality of I-LLMRec in reality.
Table 9: Performance under low-quality images.
Row Level of Image Quality Hit@5 NDCG@5
Using only image features
(a) I-LLMRec w/ low quality images 0.3704 0.2936
(b) I-LLMRec w/ high quality images 0.4316 0.3403
Using all features (Image, CF, Text)
(c) I-LLMRec w/ low quality images 0.4459 0.3597
(d) I-LLMRec w/ high quality images 0.4570 0.3711
E.9 Exploring Sensitivity to Image Quality
Throughout the paper, we use high-resolution images, i.e., high-quality item images. However, in the real-world applications, it may not always be feasible to obtain such high-quality item images, and systems may instead rely on low-quality ones. To examine this issue, we investigate how I-LLMRec performs when the low-quality images are used. Specifically, we replace high-resolution images with low-resolution versions originally provided in the metadata. We evaluate this setting on the Amazon Sport dataset under two conditions: (i) using only image features in both training and inference to isolate the effect of image quality, and (ii) using all features [^14] The LLM takes only the item images consumed by users as input, whereas the retrieval-based recommendation leverages image, text, and CF features. (image, text, and CF). As shown in Table 9, we observe that performance drops significantly when relying solely on image features (row (a) vs. (b)), while the model remained robust when other features were included (row (c) vs. (d)). These observations indicate that while I-LLMRec is sensitive to image quality when used in isolation, it can still effectively handle low-quality images by leveraging complementary information.
Table 10: Training Strategy.
Training Strategy Sport Art
@@ E.9 Exploring Sensitivity to Image Quality
End-to-End 0.4570 0.3711 0.5883 0.4839
Stage1+Stage2 0.4604 0.3747 0.5858 0.4887
E.10 Training Strategy
Throughout this paper, we optimize our model using an end-to-end training strategy to ensure ease of implementation and learning efficiency. However, in this section, we explore the impact of an alternative two-stage training strategy, where we separate the RISA and RERI modules. Specifically, for stage 1, we train the adaptor $M$ through the RISA module to initially align the visual space with the language space. For stage 2, we freeze the adaptor and train only $F_{u}^{*}$ and $F_{i}^{*}$ (*=[Img, CF, Text]) through the RERI module. As shown in Table 10, the two-stage strategy shows performance comparable to the end-to-end strategy, indicating that neither strategy offers a clear advantage over the other. Given this, we opt for the end-to-end training strategy to facilitate the ease of implementation and training efficiency.
Figure 12: Case studies of impact of figure types in images.
E.11 Case Study on the Impact of Figure Types
To study which figure types influence performance, we analyzed the item histories of users who correctly predicted the test item and selected their top 5 most frequently consumed items (positive set), as well as the top 5 items from users who failed to predict the test item correctly (negative set), using the Amazon Sport and Art datasets. We hypothesize that items in the positive set help LLMs capture user preferences, whereas items in the negative set have the opposite effect. As shown in Figure 12, we observe that in the Sport dataset, the positive set mainly consists of sports apparel (e.g., shoe, athletic shirts), while the negative set includes tactical or survival-related gear (e.g., camping pillow, lightbulb). This suggests that the functional category or intended use of an item appears to have an influence on performance. On the other hand, in the Art dataset, items in the positive set typically feature clear and easily recognizable visuals (e.g., paintbrush, yarn), whereas the negative set contains items that require interpreting small text or fine details to identify (e.g., a dark chipboard with a small label). These findings indicate that images providing clear visual cues are more beneficial for the model than those requiring detailed or text-based interpretation, instead of the functional category effect observed in the Sport dataset. In conclusion, the types of figures within images that affect performance seem to vary depending on the dataset.
Table 11: Impact of the different backbones.
@@ E.11 Case Study on the Impact of Figure Types
E.12 Impact of Different LLMs and Vision Encoders
@@
To investigate the impact of the backbone models on the recommendation performance of  I-LLMRec, we evaluated our model by replacing different backbone models. For the LLM, we used Sheared LLaMA-2.7B Xia et al. (2023) and Vicuna 1.5-7B Chiang et al. (2023), an improved version of LLaMA-7B Touvron et al. (2023a). For the vision encoder, we used CLIP [^15] https://huggingface.co/openai/clip-vit-large-patch14-336 Radford et al. (2021) and SigLIP [^16] https://huggingface.co/google/siglip-so400m-patch14-384 Zhai et al. (2023), with the advanced version of CLIP scaling with sigmoid loss. Table 11 shows the following observations: 1) Given the same vision encoder, I-LLMRec with Vicuna 1.5-7B outperforms the model with LLaMA-2.7B. This improvement is likely attributed to Vicuna 1.5-7B’s superior semantic reasoning capability, allowing it to capture user preferences more effectively. 2) Given the same LLM backbone, I-LLMRec with SigLIP performs better than the version with CLIP. This suggests that models that better capture visual features are more effective in delivering item semantics, thereby enabling the LLM to better understand the item semantics. In summary, employing more powerful LLMs and vision encoders leads to improvements in recommendation performance, as these models effectively capture item semantics and user preference.
E.13 Simulation of Scaling Item Catalog
We explore how  I-LLMRec manages the scalability of the item catalog by simulating its performance on the Amazon Phone dataset. In this simulation, we gradually increase the item catalog size from 5K to 20K (the full set). For a realistic setting, we assume that for each catalog size, there is a corresponding user base that has interacted solely with the items in that catalog. Specifically, we model the following user counts: 2.7K, 13.8K, 33.7K, and 59.3K users for catalog sizes of 5K, 10K, 15K, and 20K, respectively. This approach simulates the growing user base of expanding online services.
The performance of the models trained on each item catalog shows values of 0.468, 0.463, 0.4701, and 0.5106, indicating relatively stable performance even as the catalog size grows. However, we recognize that frequent model retraining results in increased memory and computational costs. To mitigate this, we tested using a model trained on a smaller catalog to infer for a larger catalog (e.g., a model trained on 5K items inferring for 10K, and a model trained on 10K items inferring for 15K). This approach avoids the need for periodic retraining as the catalog size grows and assess how robustly past models can perform as the item catalog grows. We observe that the performance remains robust with values of 0.370 (5K model to 10K catalog), 0.422 (10K model to 15K catalog), and 0.444 (15K model to 20K catalog), respectively, thanks to the I-LLMRec’s cold-start item recommendation ability (See Appendix E.5). These results suggest that I-LLMRec can operate efficiently without frequent retraining, significantly reducing both computational complexity and training costs. Regarding system performance, I-LLMRec achieves a throughput of 20 users per second, including the indexing of visual features1, across all catalog sizes. This assumes the use of a fast retrieval model to first retrieve relevant items from the larger pool, with I-LLMRec performing re-ranking within the top 100 candidates. The GPU memory requirements are as follows: 6.7GB to load the model, and 21GB of GPU memory during inference with a batch size of 16. In terms of storage, the model weights require 14.3GB, and the visual feature cache requires 3.4KB per item. Thus, for every 5K increase in the catalog size, an additional 17MB of storage is needed. However, as the item catalog grows, the I-LLMRec model weights replace the previous model weights, meaning that no additional storage is required. For cost per request, since LLM-based recommendation models are primarily used for re-ranking [1], the processing cost remains constant regardless of catalog size, as indexing the visual features is negligible with a complexity of O(1). The key scaling bottleneck is the periodic retraining necessary to maintain performance as the catalog expands. Using a single NVIDIA GeForce A6000 48GB GPU, an additional 3 hours of training time is required for every 5K increase in catalog size. However, we anticipate that utilizing multiple GPUs for offline training and fine-tuning models trained on previous catalogs will substantially reduce training time, compared to the current approach, which retrains the model from scratch with each catalog size increase.
Appendix F Additional Related Works
Multimodal LLM-based Recommendation. Recent studies (Zhou et al., 2025; Ye et al., 2025; Wei et al., 2024) have explored leveraging visual information in LLMs for recommendation tasks. Specifically, MSRBench (Zhou et al., 2025) investigates the use of off-the-shelf Multimodal LLMs (e.g., GPT-4V (Achiam et al., 2023)) for recommendation in various settings, while MLLM-MSR (Ye et al., 2025) and UniMP (Wei et al., 2024) fine-tune the Multimodal LLMs (Awadalla et al., 2023; Liu et al., 2023a) developed in computer vision to improve recommendation performance. Nonetheless, these studies mainly emphasize leveraging Multimodal LLMs to merely boost performance by feeding image features into LLMs, offering limited analysis of image–description overlap and lacking effective alignment strategies between visual and language spaces tailored to the recommendation context.
Appendix G Limitation and Future work
In this paper, we mainly exploit item images to capture user preferences effectively and efficiently for LLMs. A potential limitation of this approach is the possible underutilization of textual information. Specifically, even if item images are sufficient to capture user preferences by offering rich semantics, textual descriptions may still provide complementary information—such as subtle attributes or contextual nuances—that are not easily conveyed by images. However, extracting such information requires carefully disentangling text-specific cues from overlapping visual content and isolating only those parts of text that contribute meaningfully to recommendation tasks, which is a challenging and non-trivial process. Although the contribution of this information is estimated to be small according to our results in Section 4.2, combining it with image features could further enhance recommendation performance. We therefore identify this as a promising direction for future research.